OpenStudy (anonymous):

Solve for the variable. 2/3x =26 Answer 39 -39 17.333... -17.333...

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 }x =26\]

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 }x=26 \rightarrow x=26 \div \frac{ 2 }{ 3 }=26 \times \frac{ 3 }{ 2 }=\frac{ 78 }{ 2 }=39\]Therefore, x = 39. @JennWoww

OpenStudy (anonymous):

thank you @genius12 :)

OpenStudy (anonymous):

No problem.